a light ray is incident on a plane mirror . the angle deviation is 120, then at what angle does incident ray make with the mirror?

Grade:9

1 Answers

Ananya Kapoor
7 years ago
As its deviation is 120 degree. Angle i = Angle r. So i=r=120/2=60 degree. We know that angle of incident ray from mirror + angle i =90 degree. So the answer is 90-60= 30 degree